### v2

@@ Abstract
Nonlinear optical interactions in nanostructures are essential for both fundamental research and practical applications. The optical Kerr effect, a third-order nonlinear optical phenomenon, induces anisotropic changes in the refractive index of a material when exposed to high-intensity laser beams. However, the relatively modest magnitude of this effect limits its practical application in free-space nanophotonics. In this study, we achieve an enhancement of the optical Kerr effect by over three orders of magnitude using guided-mode resonance in a dielectric diffraction grating. Our work includes the design, fabrication, and measurement of the transmittance of a resonant silicon nitride (SiNx) crossed grating illuminated with short light pulses. We observe spectral resonance shifts with varying pulse power, resulting in more than fifty percent modulation of the resonance magnitude. Additionally, this study investigates the potential of the grating for dynamic pulse shaping. The presented results unveil promises in the development of next-generation all-optical devices, leveraging enhanced Kerr nonlinearities in dielectric gratings.
@@ Introduction
Since the groundbreaking experiments on second-harmonic generation (SHG) in 1961 [1] and optical wave mixing in 1962 [2], nonlinear optics has grown substantially. Contemporary research is increasingly focused on integrating nonlinear optical materials into miniaturized devices [3, 4]. This progress is driven by advancements in nanofabrication and material science, which have led to the development of novel materials, such as metal-dielectric nanocomposites [5] and epsilon-near-zero materials [6], both known for their significant nonlinear optical responses.
@@
The optical Kerr effect (OKE), a third-order nonlinear phenomenon, induces anisotropic refractive index changes under intense laser illumination [7]. This effect has been utilized in integrated optics to develop all-optical switches, logic gates, and bistable optical elements [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], and has been explored in free-space diffractive elements as well [19]. While free-space structures are suitable for large-scale parallel all-optical signal processing, they face challenges with short light-matter interaction lengths, limiting volumetric nonlinear optical effects. To address this, researchers have turned to materials with significant nonlinear properties and strong electromagnetic field confinement, achieved either through material engineering or by exploiting resonances in sub-wavelength diffractive structures such as metamaterials [4]. Although plasmonic structures enhance nonlinear signals [20], their high Ohmic losses limit effectiveness. All-dielectric nanophotonics, offering negligible losses and high damage thresholds [21, 22, 23], provides a promising alternative.
@@
Among all-dielectric structures, guided-mode resonant (GMR) gratings stand out for their high-Q resonance, strong field confinement, and tunability [24]. Initially introduced by Mashev and Popov [25, 26] and further explored by Golubenko et al. [27], GMRs have applications in laser resonators, narrow-band filters, and selective mirrors [28, 29]. While early studies focused on static devices, dynamic tuning via quadratic electro-optic effects in GMR gratings was later introduced by Sharon et al. [30], leading to further exploration of mechanical and nonlinear tuning mechanisms. Numerical studies on GMR-based optical switching and bistability [31, 17, 32, 33] and experimental studies on tunable GMRs using slow nonlinear effects in Azobenzene dye [34, 35] have expanded understanding of the potential of these structures.
@@
In this study, we present the first experimental demonstration of ultrafast nonlinear optical modulation in a guided-mode resonance (GMR) grating. Our results show over 50% modulation of the resonance magnitude across the spectrum, driven by an OKE-induced shift in the resonance peak. For this experiment, we fabricated a grating with square-shaped pillars arranged in a square lattice, utilizing low-loss SiNx material, which exhibits moderately high intrinsic self-defocusing nonlinearity within the relevant spectral range. The experimental results are validated through comprehensive numerical simulations. By employing a semi-analytical approach, we elucidate the mechanisms underlying the waveguide-enhanced nonlinear optical effects and the resulting resonance peak shift. To gain deeper insight into the origins of the GMR-enhanced OKE, we visualize the field and permittivity distributions within the structure. Additionally, we demonstrate the potential of the grating structure for all-optical pulse shaping and nonlinear material characterization.
Results and Discussion
Theory and device concept
@@
Figure 1: Geometry of the GMR grating. (a) Three-dimensional (3D) perspective of an optical pulse incident on the grating from air. One grating period is marked with the dotted rectangle. (b) Side view showing that the wave vector $\mathbf{k}_{0}$ is along the positive $z$-axis, and the electric field ($\mathbf{E}$) is polarized along $y$-axis. (c) A scanning electron micrograph (top-view) of the fabricated GMR grating.
Numerical simulations and experimental results
@@
The diffraction efficiencies of the GMR in direct transmission are plotted against the wavelength in air ($\lambda_{0}$) for two different average powers of the incident light, $P_{1}=0.5$ mW (solid lines), and $P_{1}=4.5$ mW (dashed lines). The plots in ‘Black’ color represent numerical simulation results, while the ‘Red’ lines represent experimental results in Fig. 2(b). The third-order susceptibility $\chi^{\mathrm{(3)}}$ of SiNx is measured with a custom built $z$-scan set-up. To get adequate nonlinear signal in the $z$-scan measurements, a thicker SiNx film on a fused silica substrate was used. This thicker film possesses identical optical properties to the one used in the grating fabrication, as both were synthesized under identical deposition conditions. Numerical simulations are conducted using the Fourier Modal Method (FMM) for gratings made with materials exhibiting Kerr-type nonlinearity [39]. $\chi^{\mathrm{(3)}}$ of SiNx measured with $z$-scan is used in the nonlinear FMM simulations. Moreover, we assume that the origin of the nonlinearity in SiNx is the nonlinear response of bound electrons to the applied electromagnetic field. Hence, $B^{\prime}/A^{\prime}=2$ is used in Eq. (8) [7]. The complete set of numerical simulation and experimental results is included in Section S5 of the supplementary information. We observe a reduction in the modulation depth of the resonances in the experimental results compared to the numerical simulations. This discrepancy can be attributed to the finite size and reduced spatial coherence of the beam. For the results obtained with higher powers of the laser beam, the two-photon absorption of the SiNx material also reduces the modulation depth due to non-negligible nonlinear optical losses. The two-photon absorption coefficient $\beta^{\mathrm{TPA}}$ is also measured in $z$-scan. The experimental and the fitted $z$-scan results are included in section S5 of the supplementary information. The original and the shifted resonance peak positions $\lambda_{1}$, and $\lambda_{2}$ are marked by the ‘Green’ and the ‘Blue’ circles in Fig. 2(b). Examining at $\lambda_{2}$, a relative change of over $50\%$ in the diffraction efficiency is observed due to the nonlinear optical modulation of the GMR.
@@
Electric field and permittivity distributions inside the GMR
@@
Summary and Discussion
To summarize, we present the first experimental demonstration of ultrafast nonlinear optical modulation of guided-mode resonance (GMR) in a diffraction grating. The grating, fabricated using a CMOS-compatible low-loss SiNx material, exhibits more than three orders of magnitude enhancement in the Optical Kerr Effect (OKE) compared to a bare film of the same thickness. The experimental results presented in this article reveal over 50% modulation of the resonance magnitude across the spectrum, attributed to Kerr nonlinearity-induced shifts of the resonance peak. These experimental findings are validated through comprehensive numerical simulations and a semi-analytical approach for qualitative understanding. Visualization of the field and permittivity distributions within the grating enhances our understanding of the origin of the GMR-enhanced OKE, leading to strong and tunable light-induced anisotropy.
We also demonstrate the feasibility of GMR-enhanced OKE for all-optically tunable pulse shaping. Furthermore, the excellent agreement between the Kerr nonlinearity characterized by the resonance peak (or dip) shift and that obtained from $z$-scan measurements of a bare SiNx film suggests that GMR-enhanced OKE can be exploited to devise an alternative technique for nonlinear optical material characterization. Unlike conventional methods that require focusing pulsed light on a small area, this alternative technique would allow for measurements with a collimated beam, thereby reducing the risk of material damage. Overall, this research opens new avenues for utilizing Kerr-type nonlinearities in all-dielectric gratings and paves the way for advancements in high-resolution nonlinear imaging, laser engineering, and all-optical signal processing.
@@ Methods
Acknowledgments S. B. gratefully acknowledges the insightful discussions with Emeritus Prof. (late) Jari Turunen during the initial planning stages of this work. The authors also extend their sincere thanks to Dr. Hannu Pasanen for his support in establishing the experimental setup, and to Dr. Petri Karvinen for his essential advice during the fabrication of the grating structure.
@@
  • Authors’ contributions- S.B. conceived the idea, planned the research with T.N., and designed and fabricated the grating structure. S.B. also established the experimental setup, performed the measurements, and analyzed the experimental data. T.N. supported the numerical simulations and the analysis of the experimental data. N.T. assisted in setting up the experiments and provided valuable guidance for the alignment of the laser and the OPA. T.N. and R.F. provided suggestions throughout the progress of the work. S.B. prepared the manuscript with input from T.N. S.B., T.N., and R.F. collaboratively edited the manuscript.

### v3

@@ Abstract
Nonlinear optical interactions in nanostructures are subjects of both fundamental research and practical applications. The optical Kerr effect, a third-order nonlinear optical phenomenon, induces anisotropic changes in the refractive index of a material under intense laser illumination. This effect is inherently weak, limiting its practical use in free-space nanophotonics. In this work, we demonstrate an enhancement of the optical Kerr effect by more than three orders of magnitude using guided-mode resonance in a dielectric diffraction grating with an overall thickness of less than 300 nm. Our study involves the design, fabrication, and measurement of the transmittance of a resonant silicon nitride crossed grating illuminated with short light pulses. We observe spectral resonance shifts with varying pulse power, leading to more than fifty percent modulation of the resonance magnitude and enabling transitions between ‘Off’ and ‘On’ states of transmission. We also demonstrate the capability of the grating for dynamic pulse shaping. The results presented here reveal promising directions for developing advanced all-optical devices based on free-space nanophotonics.
@@ Introduction
Since the pioneering experiments on second-harmonic generation (SHG) [1], and optical wave mixing [2], the field of nonlinear optics has grown substantially. Recent research has focused on integrating novel nonlinear optical materials into miniaturized devices for practical applications [3, 4]. This progress has been driven by advancements in nanofabrication and materials science, which have enabled the development of innovative materials, such as metal-dielectric nanocomposites [5] and epsilon-near-zero materials [6], both recognized for their remarkable nonlinear optical responses.
@@
The optical Kerr effect (OKE), a third-order nonlinear phenomenon, induces anisotropic refractive index changes in materials under intense laser illumination [7]. This effect has been harnessed in integrated optics to develop various components, including all-optical switches, logic gates, and bistable optical elements [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. OKE has also been explored in free-space diffractive elements [19]. While free-space structures are advantageous for large-scale parallel all-optical signal processing, they encounter challenges due to short light-matter interaction lengths, which limit volumetric nonlinear optical effects.
To address these limitations, researchers have concentrated on synthesizing materials with stronger nonlinear optical responses and leveraging strong electromagnetic field confinement in optically thin materials. These objectives can be achieved through advanced material engineering and by exploiting resonances in sub-wavelength diffractive structures, respectively [4]. Although plasmonic structures can enhance nonlinear signals [20], their high Ohmic losses limit their practical applications. In contrast, all-dielectric nanophotonics presents a promising alternative, offering negligible losses and high damage thresholds [21, 22, 23].
@@
Among all-dielectric structures, guided-mode resonant (GMR) gratings stand out for their high-Q resonance, strong field confinement, and tunability [24]. Initially introduced by Mashev and Popov [25, 26] and further explored by Golubenko et al. [27], GMR gratings have been proven to be useful in laser resonators, narrow-band filters, and selective mirrors [28, 29]. While early studies focused on static devices, dynamic tuning via quadratic electro-optic effects in GMR gratings was later introduced by Sharon et al. [30], leading to further exploration of mechanical and nonlinear tuning mechanisms. Numerical studies on GMR-based optical switching and bistability [31, 17, 32, 33] and experimental studies on tunable GMRs using slow nonlinear effects in Azobenzene dye [34, 35] have expanded understanding of the potential of these structures.
@@
In this study, we present the first experimental demonstration of ultrafast nonlinear optical modulation in a guided-mode resonant (GMR) grating. Our results show over 50% modulation of the resonance magnitude across the spectrum, driven by an optical Kerr effect (OKE)-induced shift in the resonance peak, which enables dynamic switching between ’On’ and ’Off’ transmission states of the grating. For this experiment, we fabricated a grating consisting of square-shaped pillars arranged in a square lattice, using low-loss SiNx material, which exhibits moderately high intrinsic self-defocusing nonlinearity within the relevant spectral range. The experimental findings are supported by comprehensive numerical simulations.
We also employ a semi-analytical approach to elucidate the mechanisms underlying the waveguide-enhanced nonlinear optical effects and the associated resonance peak shift. To further understand the origins of the GMR-enhanced OKE, we visualize the field and permittivity distributions within the structure. Additionally, we demonstrate the potential of this grating structure for all-optical pulse shaping and nonlinear material characterization.
Theory and Results
Semi-analytical model
@@
Figure 1: Geometry of the GMR grating. (a) Three-dimensional (3D) perspective of an optical pulse incident on the grating from air. One grating period is marked with the dotted rectangle. (b) Side view showing that the wave vector $\mathbf{\kappa}_{0}$ is along the positive $z$-axis, and the electric field ($\mathbf{E}$) is polarized along $y$-axis. (c) A scanning electron micrograph (top-view) of the fabricated GMR grating.
Full-wave numerical simulations and experimental results
@@
The diffraction efficiencies of the GMR in direct transmission are plotted against the wavelength in air ($\lambda_{0}$) for two different average powers of the incident light, $P_{1}=0.5$ mW (solid lines), and $P_{1}=4.5$ mW (dashed lines). The plots in ‘Black’ color represent numerical simulation results, while the ‘Red’ lines represent experimental results in Fig. 2(b). The third-order susceptibility $\chi^{\mathrm{(3)}}$ of SiNx is measured with a custom built $z$-scan set-up. To get adequate nonlinear signal in the $z$-scan measurements, a thicker SiNx film on a fused silica substrate was used. This thicker film possesses identical optical properties to the one used in the grating fabrication, as both were synthesized under identical deposition conditions. Numerical simulations are conducted using the Fourier Modal Method (FMM) for gratings made with materials exhibiting Kerr-type nonlinearity [39]. $\chi^{\mathrm{(3)}}$ of SiNx measured with $z$-scan is used in the nonlinear FMM simulations. Moreover, we assume that the origin of the nonlinearity in SiNx is the nonlinear response of bound electrons to the applied electromagnetic field. Hence, $B^{\prime}/A^{\prime}=2$ is used in Eq. (8) [7]. The complete set of numerical simulation and experimental results is included in Section S5 of the supplementary information. We observe a reduction in the modulation depth of the resonances in the experimental results compared to the numerical simulations. This discrepancy can be attributed to the finite size and reduced spatial coherence of the beam. For the results obtained with higher powers of the laser beam, the two-photon absorption of the SiNx material also reduces the modulation depth due to non-negligible nonlinear optical losses. The two-photon absorption coefficient $\beta^{\mathrm{TPA}}$ is also measured in $z$-scan. The experimental and the fitted $z$-scan results are included in section S5 of the supplementary information. The original and the shifted resonance peak positions $\lambda_{1}$, and $\lambda_{2}$ are marked by the ‘Green’ and the ‘Blue’ circles in Fig. 2(b). Examining at $\lambda_{2}$, a relative change of over 50% in diffraction efficiency is observed due to the nonlinear optical modulation of the GMR, effectively switching the GMR grating from a transmission ‘On’ to ‘Off’ state.
@@
Electric field and permittivity distributions
@@
Summary
In summary, we present the first experimental demonstration of ultrafast nonlinear optical modulation of guided-mode resonance in a diffraction grating. The grating, fabricated using a CMOS-compatible, low-loss SiNx material, shows an enhancement in the Optical Kerr Effect (OKE) by more than three orders of magnitude compared to a bare film of the same thickness. Our experimental results reveal over 50% modulation of the resonance magnitude across the spectrum, attributed to Kerr nonlinearity-induced shifts in the resonance peak, enabling all-optical switching between ’On’ and ’Off’ transmission states. These findings are validated through comprehensive numerical simulations and a semi-analytical approach for qualitative understanding. Visualization of the field and permittivity distributions within the grating further clarifies the origin of the GMR-enhanced OKE, resulting in strong and tunable light-induced anisotropy. Additionally, we demonstrate the potential of GMR-enhanced OKE for all-optically tunable pulse shaping.
Discussion
We observed strong agreement between the magnitude of the optical Kerr effect, indicated by the spectral shift of the resonance dip, and the results from $z$-scan measurements of a bare SiNx film. This finding highlights the potential of using guided-mode resonance (GMR)-enhanced Optical Kerr Effect (OKE) as an alternative method for nonlinear optical material characterization. Unlike traditional methods, such as the $z$-scan technique that requires tightly focused pulsed light, which can cause localized heating or damage, the GMR-enhanced OKE approach allows for measurements with a collimated beam, significantly reducing the risk of material damage.
This research opens new avenues for exploiting Kerr-type nonlinearities in all-dielectric gratings, paving the way for advancements in high-resolution nonlinear imaging, where enhanced responses improve contrast and sensitivity, and in laser engineering, where precise light-matter interaction is key. The findings also have important implications for all-optical signal processing, which demands fast, tunable, low-power components. The ability to dynamically modulate light using all-dielectric materials without high losses offers a promising direction for developing next-generation photonic devices and systems.
@@ Methods
Acknowledgments S. B. gratefully acknowledges the insightful discussions with the late Emeritus Prof. Jari Turunen, which led to the concept behind this project. The authors also extend their sincere thanks to Dr. Hannu Pasanen for his support in establishing the experimental setup, and to Dr. Petri Karvinen for his essential advice during the fabrication of the grating structure.
@@
  • Authors’ contributions- S.B. conceived the idea behind this work, planned the research activities jointly with T.N. S.B. designed and fabricated the grating structure, established the experimental setup, performed the measurements, and analyzed the experimental data. T.N. supported the numerical simulations and the analysis of the experimental data. N.T. assisted in setting up the experiments and provided valuable guidance for the alignment of the laser and the OPA. T.N. and R.F. provided suggestions throughout the progress of the work. S.B. prepared the manuscript with inputs from T.N. S.B., T.N., and R.F. collaboratively edited the manuscript.